OpenStudy (anonymous):

Find all solutions to 5sinx+2=0 on [0,2pi], rounding to three decimal places...

OpenStudy (anonymous):

i have 2.5 minutes left on this practice test to answer!

OpenStudy (anonymous):

5sinx+2=0 5 sinx =-2 sinx =-2/5 Now, use calculator to find x= arcsin(-2/5)

OpenStudy (anonymous):

can u?

OpenStudy (arhin):

5sinx+2=0 5sinx=-2 sinx= -2/5 now,locate the quadrant where sin is negative try it and u will get the ans right